# Global monopole in a Ricci-coupled Kalb-Ramond bumblebee gravity

### v2

@@ Abstract
In this paper, we investigate black hole solutions in Einstein-Kalb-Ramond (EKR) bumblebee gravity sourced by a global monopole characterized by the charge $\eta$. This modified theory of gravity possesses the notable feature of incorporating local Lorentz symmetry breaking (LSB) via a spontaneous symmetry-breaking mechanism. We solve the field equations for a static and spherically symmetric metric with the Kalb-Ramond (KR) field fixed at its VEV, thereby obtaining new black hole solutions. These solutions simultaneously exhibit the LSB effects, codified by the $\gamma$ parameter, and the global monopole effects, codified by the charge $\eta$. Next, we study the impact of the global monopole and LSB corrections on two classical tests, namely, the advance of Mercury’s perihelion and the light deflection. Furthermore, we compute the Hawking temperature, black hole shadows, and greybody factors. Ultimately, we estimate an upper bound for $\eta$ by comparing the theoretical results provided by the EKR model with observational data from the advance of Mercury’s perihelion and light deflection.
@@ III Black hole solution with global monopole
$\mathcal{L}_{m}=\frac{1}{2}\partial_{\mu}\psi^{a}\partial^{\mu}\psi^{a}-\frac{\chi}{4}(\psi^{a}\psi^{a}-\eta^{2})^{2}.$ (10)
@@
$T^{m}_{\mu\nu}=\mathrm{diag}\bigg{(}F(r)\,\frac{\eta^{2}}{r^{2}},\frac{1}{F(r)}\,\frac{\eta^{2}}{r^{2}},0,0\bigg{)}.$ (12)
@@
This configuration leads to the constant norm $b_{\mu\nu}b^{\mu\nu}=|b|^{2}$. It is easy to verify that the field strength $H_{\lambda\mu\nu}$ identically vanishes. Then, the equations of motion for $B_{\mu\nu}$ are satisfied. Initially, we will focus on solution without cosmological constant. In this case, we adopt a quartic potential, i.e., $V(X)=\lambda X^{2}$, where $\lambda$ is a coupling constant, and $X=B_{\mu\nu}B^{\mu\nu}-b^{2}$. For this potential, the configuration (13) implies $V=V^{\prime}=0$.
@@ IV.2 Light deflection and gravitational lensing
As one can observe, the first term provides the angular deflection when $M=0$, leading to a non-zero light deflection. On the other hand, the second term represents the angular deflection for the black hole with mass $M$. One still notes that both terms bring contributions of Lorentz symmetry violation. Having the relationship of the deflection angle and the impact parameter, we are able to find the classical differential scattering section at the limit of small through $\delta\phi$ the following equation Anacleto:2022shk
@@ IV.2.2 Black hole shadow
It is straightforward to notice that the LV parameter modifies the sizes of the black hole shadows. The Fig. (1) depicts the shadow size, showing that the size decreases with the LV parameter and increases with the influence of the monopole term.
@@
Figure 1: The circles in the figure illustrate shadows formed by varying the parameters $\gamma$ for $\kappa\eta^{2}=0.1$ (a) and $\eta$ for $\gamma=0.1$ (b).
@@ VI Greybody factor
The GF depends on both the LV parameter and on the global monopole term. We depict the behavior of GF in Fig.(2). As we can see, when the frequency is at a lower level, we observe that radiation is reflected totally and there is no transmission. The tunneling effect occurs as the frequency increases so that the radiation can pass through the potential barrier. Besides, the greybody factor decreases with larger values of the LV parameter, while it almost does not change as we modify the global monopole term. In this case, we physically interpret this as follows: when the parameter $\gamma$ increases, the height of the potential barrier also increases, leading to a lower probability of transmission to the incident wave.

### v3

@@ Abstract
In this paper, we investigate black hole solutions in Einstein-Kalb-Ramond (EKR) bumblebee gravity sourced by a global monopole characterized by the charge $\eta$. This modified theory of gravity possesses the notable feature of incorporating local Lorentz symmetry breaking (LSB) via a spontaneous symmetry-breaking mechanism. We solve the field equations for a static and spherically symmetric metric with the Kalb-Ramond (KR) field fixed at its VEV, thereby obtaining new black hole solutions. These solutions simultaneously exhibit the LSB effects, codified by the $\gamma$ parameter, and the global monopole effects, codified by the charge $\eta$. Next, we study the impact of the global monopole and LSB corrections on two classical tests, namely, the advance of Mercury’s perihelion and the light deflection. Furthermore, we compute the Hawking temperature, black hole shadows, and greybody factors. Ultimately, we investigate the processes of scattering and absorption of a massless scalar field by employing the partial wave method.
@@ III Black hole solution with global monopole
$\mathcal{L}_{m}=-\frac{1}{2}\partial_{\mu}\psi^{a}\partial^{\mu}\psi^{a}-\frac{\chi}{4}(\psi^{a}\psi^{a}-\eta^{2})^{2}.$ (10)
@@
$T^{m}_{\mu\nu}=\mathrm{diag}\bigg{(}F(r)\,\frac{\eta^{2}}{r^{2}},-\frac{1}{F(r)}\,\frac{\eta^{2}}{r^{2}},0,0\bigg{)}.$ (12)
@@
This configuration leads to the constant norm $b_{\mu\nu}b^{\mu\nu}=-|b|^{2}$, and $X=B^{\mu\nu}B_{\mu\nu}+b^{2}$. It is easy to verify that the field strength $H_{\lambda\mu\nu}$ identically vanishes. Then, the equations of motion for $B_{\mu\nu}$ are satisfied. Initially, we will focus on solution without cosmological constant. In this case, we adopt a quartic potential, i.e., $V(X)=\lambda X^{2}$, where $\lambda$ is a coupling constant, and $X=B_{\mu\nu}B^{\mu\nu}-b^{2}$. For this potential, the configuration (13) implies $V=V^{\prime}=0$.
@@ IV.2 Light deflection and gravitational lensing
As one can observe, the first term provides the angular deflection when $M=0$, leading to a non-zero light deflection. On the other hand, the second term represents the angular deflection for the black hole with mass $M$. One still notes that both terms bring contributions of Lorentz symmetry violation. Having the relationship of the deflection angle and the impact parameter, we are able to find the classical differential scattering section at the limit of small $\delta\phi$ through the following equation Anacleto:2022shk
@@ IV.2.2 Black hole shadow
It is straightforward to notice that the LV parameter modifies the sizes of the black hole shadows. The Fig. (1) depicts the shadow size, showing that the size decreases with the LV parameter and increases with the influence of the global monopole term. Additionally, we depict the shadow radius versus the parameters $\gamma$ and $\eta$ in Fig. (2).
@@
Figure 1: The circles in the figure illustrate shadows formed by varying the parameters $\gamma$ for $\kappa\eta^{2}=0.1$ (a) and $\eta$ for $\gamma=0.1$ (b). We have set $M=0.5$.
Refer to caption Refer to caption
(a)                                                                         (b)
Figure 2: Shadow radius versus parameter the parameter $\gamma$ (a) and versus the parameter $\eta$ (b). We have set $M=0.5$.
We can use observational data from Event Horizon Telescope (EHT) related to black hole shadow of Sagittarius A* (Sgr A*), a supermassive black hole at the center of the Milky Way galaxy, to estimate bounds for the parameters $\gamma$ and $\eta$. Firstly, let us define the radius of the observed shadow of Sgr A* as follows
$R_{SgrA*}=\frac{D\,\Theta_{s}}{2M_{SgrA*}},$ (82)
where $\Theta_{s}$ represents the angular size of the black hole shadow of Sgr A* galactic center reported by EHT, $M_{SgrA*}$ is the mass of Sgr A* and $D$ is the distance to Sgr A*. According to EHT data based on estimative from the Keck and Very Large Telescope Interferometer (VLTI) instruments, one has $\Theta_{s}=(48.7\pm 7)\,\mu as$, $M_{SgrA*}=4\times 10^{6}M_{\odot}$ and $D=8\,kpc$ EventHorizonTelescope:2022wkp ; EventHorizonTelescope:2022exc ; EventHorizonTelescope:2022xqj . With this data in hands, we obtain the shadow radius of Sgr A* in mass units given by
$R_{SgrA^{*}}=(5.0\,\pm 0.5)M_{SgrA*}.$ (83)
In order to make a comparison between the observational and theoretical values, let us consider some values for the parameters $\gamma$ and $\eta$. Firstly, from the shadow radius (81) with $\gamma=\eta=0$, we get the theoretical result
$r_{s}=5.196M_{SgrA*}.$ (84)
If we set $\gamma=0.10$ and $\kappa\eta^{2}=0.15$, we find the following result
$r_{s}=5.50M_{SgrA*},$ (85)
which is the upper limit of the shadow radius of Sgr A*. On the other hand, by setting $\gamma=0.25$ and $\kappa\eta^{2}=0.15$, one finds
$r_{s}=4.58M_{SgrA*},$ (86)
which is the lower limit of the shadow radius of Sgr A*. In general, to agree with the observational data of the shadow radius of Sgr A*, the parameters $\gamma$ and $\eta$ must satisfy the following relation
$0.867\lesssim\frac{1-\gamma}{1-\kappa\eta^{2}}\lesssim 1.059,$ (87)
which imposes bounds on the parameters $\gamma$ and $\kappa\eta^{2}$.
@@ VI Greybody factor
The GF depends on both the LV parameter and on the global monopole term. We depict the behavior of GF in Fig.(3). As we can see, when the frequency is at a lower level, we observe that radiation is reflected totally and there is no transmission. The tunneling effect occurs as the frequency increases so that the radiation can pass through the potential barrier. Besides, the greybody factor decreases with larger values of the LV parameter, while it slightly increase as we modify the global monopole term. In this case, we physically interpret this as follows: when the parameter $\gamma$ increases, the height of the potential barrier also increases, leading to a lower probability of transmission to the incident wave.